Calculate the [OH−] and the pH of a solution with an [H+]=4.0×10−13 M at 25 °C .

[OH−]= 

pH=


Calculate the [H+] and the pH of a solution with an [OH−]=9.8×10−13 M at 25 °C .

[H+]= 

pH=



Calculate the [H+] and the [OH−] of a solution with a pH=3.52 at 25 °C .

[H+]= 

[OH−]= 



1
Expert's answer
2022-04-26T06:14:02-0400

pH+pOH=14, pH=-lg[H+],

1)pH=-lg(4.0×10^−13)=12.4, pOH=14-12.4=1.6;

2)pOH=-lg(9.8*10^-13)=12, pH=2. [H+]= 10^-2= 0.01;

3)pOH=14-3.52=10.48, [H+]=10^-3.52=3*10-4 mol/l  [OH−]=10^-10.48 = 3.3*10^-11 mol/l